(1+cosx^2+2cosx)/(1+cosx^2-2cosx)
Answer by addingup(3677)

1+cos^2(x)+2cos(x))/(1+cos^2(x)-2cos(x)) = 0
Simplify, 2cos(x)/-2cos(x) eliminate each other but we must retain the -:
(1+cos(x))^2/(cos(x)-1)^2 = 0
Substitute y = cos(x):
(y+1)^2/(y-1)^2 = 0
Multiply both sides by (y-1)^2:
(y+1)^2 = 0
Take the square root of both sides:
y+1 = 0
Subtract 1 from both sides:
y = -1
Substitute back for y = cos(x):
cos(x) = -1
Take the inverse cosine of both sides and you have your answer:
x = pi+2pi(n)